OpenStudy (anonymous):

confused on percent problems ex: what percent of 126 is 22? PLEASEE HELPP

jimthompson5910 (jim_thompson5910):

what percent of 126 is 22 (x/100)*126 = 22 126x = 22*100 126x = 2200 I'll let you finish

OpenStudy (anonymous):

17.5?

jimthompson5910 (jim_thompson5910):

I'm getting 17.46031746, but if you round it to one decimal place, you will get 17.5 % round it to 2 decimal places to get 17.46%

OpenStudy (anonymous):

ok but i am not 100% good at this i just don't know how to do it for some

OpenStudy (anonymous):

what is ur strategy

jimthompson5910 (jim_thompson5910):

well the "what percent" part is x/100 since x% = x/100

jimthompson5910 (jim_thompson5910):

of means multiply so "what percent of 126" means (x/100)*126

jimthompson5910 (jim_thompson5910):

is means = so all together "what percent of 126 is 22" means "(x/100)*126 = 22"

OpenStudy (anonymous):

THNAKKS

OpenStudy (anonymous):

WAIT SO 75% of what # is 36...27???/

OpenStudy (phi):

75% of X (the unknown number) is 36 75/100 * X = 32 you could write this as 0.75 X =32 and divide both sides by 0.75 to get X = 32/.75 or you could simplify 75/100 to 3/4 (divide top and bottom by 25) \[ \frac{3}{4} X = 32 \] multiply both sides by 4/3 to get X = 32 * 4/3

OpenStudy (phi):

you should also think: 75% of a number is smaller than the number when you say 75% of X is 36 you know the number has to be bigger than 36

OpenStudy (anonymous):

no i actually think its 48 cuz you would do the proportion x/36=75/100

OpenStudy (phi):

? all of those ways give 48

OpenStudy (phi):

proportions are a good way to do these problems, as long as you can set up the correct proportion.